IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

  • continued work on ippToPsps & parallel dvo. all of the steps of the process now work with the parallel dvo, and I've cleaned up / simplified a bit the user interactions.

Serge Chastel

  • IPP:
    • Found out why MOPS data were not published in new format
  • MOPS:
    • Synthetics:
      • What I've implemented was "working" on Monday but I need now to test it in the IPP execution context
    • NEO cluster: IPP setup
      • Tool to register fake exposures in the IPP database (will likely be useful for other cams)
      • Problems with bash (instead of csh): fixed;
      • Dependency of chiptool on nebulous: fixed;
    • Installed recent BLAS, LAPACK, numpy, scipy, pyfits (for ipp account only) on the NEO cluster

Heather Flewelling

  • PS-Flu for 3 days.
  • czar 1 day
  • worked with serge and bryce about megacam and suprimecam processing on mops cluster
  • MD09 dvodb built except for staticsky
  • continuing to build minidvodbs for pending PSPS ingest. I need to update the wiki to reflect what's done, but the first part to ingest (kepler) is done.
  • talks with gene about ipptopsps and parallel dvo. I Need to implement parallel dvo stuff into addstar pantasks
  • worked with JT on MD stuff

Mark Huber

  • md.gr0 -- stack photometry + calibration still being cleaned up and verified. Two fields (03,05) still outstanding waiting on data shuffle disk permissions.

Bill Sweeney

  • Prototyped adding an auxiliary mask input to the chip background stage which restores the background model subtracted from the chip images. This has been requested for the M31 project. The auxiliary mask will be provided by MPG to implement more aggressive masking of bad pixels. It will only be used for the M31 data. Ran into some interesting problems with the sample mask files provided. The IPP's extra special camera configuration system failed to work with some of them. I eventually figured out what that the problem was an unexpected keyword value that caused the configuration system to choose the wrong format file. Processed the 20 sample exposures through the chip background stage, but pswarp wasn't happy with the results. Further work is required.
  • continued monitoring the LAP static sky processing. All of the stacks built prior to December have finished except for 777 from the region near the galacic center. These are now running (very slowly). Should be able to start working on the new stack in the coming week.
  • Compared the skycal results prior to and after a change made to getstar last week and didn't find any changes in the results.
  • two very busy days as processing czar.
  • out much of the week due to illness

Chris Waters

  • Finished implementing FPA-level background continuity in ppImage, with appropriate changes to camera_exp.pl to allow this to work as part of the camera stage.
  • Multiple test diffs for MOPS.
  • Began work on PATTERN.MEDIAN, the new pattern correction to remove row-to-row bias offsets.